# Differentially Private Matchings

### v1

@@ 13       Output solution, estimate
(i) follow directly from Corollary 1.3 and the fact that we require each call to Algorithm 2 to satisfy ($\nicefrac{{\varepsilon}}{{3c}}$)-privacy for $c=a\log(n)/\eta$.
@@
7.2.1 Extension to Node-Differential Privacy
We now describe the necessary modifications to adapt our edge DP continual release algorithm to the challenging node DP setting and provide the sketch pseudocode in Algorithm 7. The main task is to implement the arboricity matching sparsifier (Theorem 6.2) in an arbitrary edge-order stream. Note that there is a natural total order of vertex pairs given by the arrival order of edges. Thus we can simply discard edges where one of its endpoints has already seen more than some threshold $\Lambda$ incident edges (Algorithm 7 of Algorithm 7).
Let the sparsified graph $H$ be obtained from the input graph $G$ as described above. Write $c=O(\log(n)/\eta)$ be the maximum SVT budget for “above” queries (Algorithm 5). Our algorithm proceeds as following for each edge: Similar to Algorithm 6 of Algorithm 6, we use an SVT instance to check when to update our solution. Since the maximum matching size sensitivity is 1 for node-neighboring graphs, this step requires no change from the edge-DP algorithm. At timestamps where we must update the solution, we run a static $O(\nicefrac{{\varepsilon}}{{c\Lambda}})$-edge DP implicit matching algorithm (Algorithm 2) on the sparsified graph $H$. This is similar to Algorithm 6 of Algorithm 7 except we need to execute the underlying algorithm with a smaller privacy parameter. Finally, we also estimate the size of the largest matching contained within the current implicit solution (Lemma 6.4). Once again, this step is unchanged from Algorithm 6 of Algorithm 6. In particular, we note that we only use the sparsified graph $H$ for updating the solution and not for estimating the current maximum matching size nor for estimating the largest matching contained within the current solution.
@@ 10       Output $\operatorname{solution},\operatorname{estimate}$
The privacy proof is identical that in the proof of Theorem 1.9. The utility guarantees follow similarly, with the exception that we use the approximation guarantees of Theorem 6.3 rather than Corollary 1.3. ∎
7.2.2 Removing the Assumption on Public Bound
Finally, we describe the modifications from Algorithm 7 to remove the assumption on a public bound $\tilde{\alpha}$. First, we compute $O(\log(n))$ sparsified graphs $H_{k}$ corresponding to setting the public parameter $\tilde{\alpha}=2^{k},k=1,2,\dots,\lceil\log_{2}(n)\rceil$. The SVT to estimate the current maximum matching size remains the same (we do not compute an estimate for each $k$). Next, we compute $O(\log(n))$ implicit $b$-matchings, one for each sparsified graph $H_{k}$. We also compute an estimate of the largest contained matching for each implicit solution. Finally, note that at each time stamp, we either do nothing or update each of the $O(\log(n))$ solutions simultaneously. After such an update, we can choose the implicit solution corresponding to the smallest value of $k$ such that its estimated size is at least $\frac{1}{2+\eta}(1+\eta)^{j_{t}}-O\left(\frac{\log^{3}(n)}{\eta\varepsilon}\right)$-approximate matching.
With this, we are ready to sketch the guarantees of this modified algorithm.
Sketch Proof of Theorem 1.10.
The privacy guarantee follows from simple composition as choosing the best solution is postprocessing.
In order to show the approximation guarantees, we first show that after each solution update, one solution, estimate pair that satisfies the guarantees exists and is selected among the $O(\log(n))$ parallel instances. Let $k^{\star}$ be such that $2^{k^{\star}-1}<\alpha\leq 2^{k^{\star}}$. the $k^{\star}$-th solution certainly satisfies the desired guarantees. Remark that $(1+\eta)^{j_{t}}$ is a $\left(1+\eta,O\left(\frac{\log^{3}(n)}{\eta\varepsilon}\right)\right)$-approximate estimate of the current maximum matching size. Moreover, the estimate of the largest matching contained within the implicit solution is a $\left(1,\frac{\log^{3}(n)}{\eta\varepsilon}\right)$-approximate estimate. Thus the approximation guarantees certainly hold after each simultaneous solution update.
In between solution updates, we incur at most $(1+\eta)$-multiplicative and $O\left(\frac{\log^{3}(n)}{\eta\varepsilon}\right)$-additive error. Hence the approximation guarantees extend to all time stamps. ∎
7.3 Adjacency-List Order Streams
In this section, we give $\varepsilon$-DP algorithms for implicit matching in the continual release model with adjacency-list order streams. That is, the transcript of outputs from the release at every timestamp is $\varepsilon$-DP. The adjacency-list order stream ensures each node that arrives in this model will be followed by its edges where the edges arrive in an arbitrary order. Our algorithm is a straightforward implementation of Algorithm 2 in the continual release model.
@@ 37             $K\leftarrow K\cup\{e_{i}\}$
Privacy and Utility Guarantees
Below, we give our final theorem about the privacy and utility guarantees of our continual release algorithm. See 1.11
@@
We first prove that our continual release algorithm is $\varepsilon$-DP on the vector of outputs. First, our algorithm only outputs a new output each time a node update arrives. For every edge update, the algorithm outputs the same outputs as the last time the algorithm outputted a new output for a node update. Algorithm 8 implements all of the local randomizers used in Algorithm 2 in the following way. The order of the nodes that propose is given by the order of the node updates. Coins for edges are flipped in the same way as in Algorithm 2, the threshold for the matching condition is determined via an identical procedure, and finally, the proposal process is identical to Algorithm 2. Hence, the same set of local randomizers can be implemented as in Algorithm 2 and the continual release algorithm is $\varepsilon$-DP via concurrent composition. The approximation proof also follows from the approximation guarantee for Algorithm 2 since the proposal procedure is identical except for the additive error. In the continual release model, there is an additive error of at most $1$ since for every edge update after a new node update, the node could be matched to an initial new edge update but is not matched until the final edge update is shown for that node. Since each node contributes at most $1$ to a matching, the additive error is at most $1$ for each update. ∎
8 Improved Node-Private Bipartite Matching
The main difference in our algorithm is in implementing the counters. In the private algorithm of [hsu2014private], each of these counters need to be accurate at each of the $nT$ iterations. This causes significant privacy loss. Our algorithm has a different implementation (of the same basic algorithm), where we use the fact that only the counts at the end of each of the $T$ iterations are important. Moreover, our implementation does not need the counts themselves, but just needs to check whether they are above the supply $s$. This allows us to use the Multidimensional AboveThreshold Mechanism (Algorithm 1) to get further improvement.
3 for $t\leftarrow 1$ to $T$ do
4       foreach item $u\in V_{L}$ do
6
      // Each bidder $v$ proposes to the neighbor $u$ with lowest price
7       foreach bidder $v\in V_{R}$ do
8             if $v$ is not matched then
10                   if $p_{u}\leq 1$ then
11                         Denote bidder $v$ as (temporarily) matched with item $u$ for timestep $t$
12                         $c_{u}\leftarrow c_{u}+1$
13
14
@@ 12                         $c_{u}\leftarrow c_{u}+1$
16      foreach item $u\in V_{L}$ do
17             Release $p_{u}$
19                   $p_{u}\leftarrow p_{u}+\eta$
@@
22
23       foreach item $u\in V_{L}$ do
24             Release startu
27             while $M_{u}+\mathrm{Lap}(4/\varepsilon^{\prime})\geq\tilde{s}_{u}$ do
28                   if a bidder $v$ matched with item $u$ at timestep $\text{start}_{u}$ then
29                         Mark bidder $v$ as unmatched
30
32
33
      // Terminate early if the algorithm is close to convergence
34       Let $c_{0}$ denote the number of bids made in this iteration
36             Terminate the algorithm
37
38
8.1 Privacy Proof

### v2

@@ 8.1.2 Utility Proof
(i) follow directly from Theorem 1.3 and the fact that we require each call to Algorithm 2 to satisfy ($\nicefrac{{\varepsilon}}{{3c}}$)-privacy for $c=a\log(n)/\eta$.
@@
8.2 Proof of Theorem 1.12
8.2.1 Extension to Node-Differential Privacy with Public Arboricity Bound
We now describe the necessary modifications to adapt our edge DP continual release algorithm to the challenging node DP setting. Similary to Section 6.2.1, we first present an algorithm assuming a public bound $\tilde{\alpha}$ on the arboricity.
Theorem 8.3.
Fix $\eta\in(0,1]$. Given a public bound $\tilde{\alpha}$ on the arboricity $\alpha$ of the input graph, Algorithm 7 is an $\varepsilon$-node DP algorithm in the arbitrary order continual release model that outputs implicit $b$-matchings. Moreover, with probability at least $1-1/\operatorname{poly}(n)$, (i) $b=O\left(\frac{\tilde{\alpha}\log^{2}(n)}{\eta^{2}\varepsilon}\right)$and (ii) each implicit solution contains a $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate maximum matching.
The main task is to implement the arboricity matching sparsifier (Theorem 6.2) in an arbitrary edge-order stream. Note that there is a natural total order of vertex pairs given by the arrival order of edges. Thus we can simply discard edges where one of its endpoints has already seen more than some threshold $\Lambda$ incident edges (Algorithm 7 of Algorithm 7).
Let the sparsified graph $H$ be obtained from the input graph $G$ as described above. Write $c=O(\log(n)/\eta)$ be the maximum SVT budget for “above” queries (Algorithm 10). Our algorithm proceeds as following for each edge: Similar to Algorithm 6 of Algorithm 6, we use an SVT instance to check when to update our solution. Since the maximum matching size sensitivity is 1 for node-neighboring graphs, this step requires no change from the edge-DP algorithm. At timestamps where we must update the solution, we run a static $O(\nicefrac{{\varepsilon}}{{c\Lambda}})$-edge DP implicit matching algorithm (Algorithm 2) on the sparsified graph $H$. This is similar to Algorithm 6 of Algorithm 7 except we need to execute the underlying algorithm with a smaller privacy parameter. Finally, we also estimate the size of the current maximum matching. Once again, this step is unchanged from Algorithm 6 of Algorithm 6. In particular, we note that we only use the sparsified graph $H$ for updating the solution and not for estimating the current maximum matching size.
@@ 10       Output $\operatorname{solution},\operatorname{estimate}$
The privacy proof is identical that in the proof of Theorem 1.11. The utility guarantees follow similarly, with the exception that we use the approximation guarantees of Theorem 6.3 rather than Theorem 1.3. ∎
8.2.2 Removing the Assumption of Public Bound
Finally, we describe the modifications from Algorithm 7 to remove the assumption on a public bound $\tilde{\alpha}$. First, we compute $O(\log(n))$ sparsified graphs $H_{k}$ corresponding to setting the public parameter $\tilde{\alpha}=2^{k},k=1,2,\dots,\lceil\log_{2}(n)\rceil$. The SVT to estimate the current maximum matching size remains the same (we do not compute an estimate for each $k$). Next, we compute $O(\log(n))$ implicit $b$-matchings, one for each sparsified graph $H_{k}$. We also privately compute an upper bound $\tilde{\alpha}_{t}$ of the current arboricity of $G_{t}$ and choose the implicit solution with index $k^{\star}$ such that $2^{k^{\star}-1}<\tilde{\alpha}_{t}\leq 2^{k^{\star}}$. Accounting for the error in estimating the arboricity yields a proof of Theorem 1.12.
9 Adjacency-List Order Continual Release Implicit Matchings
In the adjacency-list order insertion streams, nodes arrive in an arbitrary order, and once a node arrives, all edges adjacent to the node arrive in an arbitrary order. Remark that in the adjacency-list order stream, all edges arrive twice, once per endpoint in the stream.
In this section, we give $\varepsilon$-DP algorithms for implicit matching in the continual release model with adjacency-list order streams. Our results for another stream type (arbitrary edge-order streams) are presented in Section 8. The adjacency-list order stream ensures each node that arrives in this model will be followed by its edges where the edges arrive in an arbitrary order. Our algorithm is a straightforward implementation of Algorithm 2 in the continual release model.
@@ 9 Adjacency-List Order Continual Release Implicit Matchings
Below, we restate our final result (Theorem 1.13) about the privacy and utility guarantees of our continual release algorithm. See 1.13
@@ 37             $K\leftarrow K\cup\{e_{i}\}$
9.1 Proof of Theorem 1.13
@@
We first prove that our continual release algorithm is $\varepsilon$-DP on the vector of outputs. First, our algorithm only outputs a new output each time a node update arrives. For every edge update, the algorithm outputs the same outputs as the last time the algorithm outputted a new output for a node update. Algorithm 8 implements all of the local randomizers used in Algorithm 2 in the following way. The order of the nodes that propose is given by the order of the node updates. Coins for edges are flipped in the same way as in Algorithm 2, the threshold for the matching condition is determined via an identical procedure, and finally, the proposal process is identical to Algorithm 2. Hence, the same set of local randomizers can be implemented as in Algorithm 2 and the continual release algorithm is $\varepsilon$-DP via concurrent composition.
The approximation proof also follows from the approximation guarantee for Algorithm 2 since the proposal procedure is identical except for the additive error. In the continual release model, there is an additive error of at most $1$ since for every edge update after a new node update, the node could be matched to an initial new edge update but is not matched until the final edge update is shown for that node. Since each node contributes at most $1$ to a matching, the additive error is at most $1$ for each update. ∎
10 Private Implicit $b$-Matchings
In this section, we extend our all of our results to the $(2+\eta)$-approximate $b$-matchings, including our node-DP results, which use a novel $b$-matching arboricity sparsifier. The key difference between the results in this section and our previous result is that our approximation bounds are given in terms of the maximum $b$-matching given large enough $b$. In particular, we satisfy the $b$-matching constraints (that our matching has degree less than $b$ for all vertices), with high probability, and output a $(2+O(\eta))$-maximum $b$-matching). Thus, our approximation is no longer a bicriteria approximation in this setting.
To obtain our $b$-matching results, we prove a stronger version of our matching result. Namely, we show that we can guarantee that our implicit output is at least the size of a $(2+\eta)$-approximate maximum $b^{\prime}$-matching for some $b^{\prime}<b$. In particular, $b$ has to be an additive factor of $\Omega\left(\frac{\ln(n)}{{\eta^{2}}\varepsilon}\right)$ greater than $b^{\prime}$ to ensure the returned solution is comparable to a $2$-approximate $b^{\prime}$-matching. Furthermore, if $b=\Omega\left(\frac{\ln(n)}{\eta^{3}\varepsilon}\right)$, then we get a $(2+\eta)$-approximate maximum $b$-matching.
10.1 $\varepsilon$-LEDP $b$-Matchings
Our revised algorithm is the same algorithm as Algorithm 2 except with a different lower bound on $b$ depending on what $b^{\prime}$-maximal matching we want to guarantee in our returned solution. This is our only change to the algorithm. When our lower bound on $b$ contains a factor of $1/\eta^{2}$, then we obtain a $(2+\eta)$-approximate maximum $b$-matching. We restate our modified algorithm below with the modified constants highlighted in blue. The constants do not affect our privacy proof; hence, we only give an updated version of our utility analysis following our pseudocode in Algorithm 2. Since the proof and algorithm only require minor changes for the generalization, we replicate our previous algorithm and proof and highlight our changes in blue.
Input: Graph $G=(V,E)$, privacy parameter $\varepsilon>0$, $b^{\prime}>0$ parameter for $b^{\prime}$-matching, matching parameter $b\geq{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}}$, constant $\eta\in(0,1)$, and constant $c\geq 3$
1 $\varepsilon^{\prime}\leftarrow\frac{\varepsilon}{2(1+2\eta)/\eta}$
4
5for each node $u\in V$ do
7       $M(u)\leftarrow\varnothing$
8       $A(u)\leftarrow\infty$
9
10for iteration $i=1$ to $n$, let $v=v_{i}$ do
12             $\nu_{i}(u)\leftarrow\text{Lap}(8/\varepsilon^{\prime})$
13             if $|M(u)|+\nu_{i}(u)\geq\tilde{b}(u)$  then
14                   $u$ releases $A(u)\leftarrow i$
@@
16
17       if $A(v)>i$ ($v$ has not satisfied its matching condition)  then
18             for subgraph index $r=0,\ldots,\lceil\log_{1+\eta}(n)\rceil$  do
20                   $|\widetilde{W}_{r}(v)|=|W_{r}(v)|+\text{Lap}(2/\varepsilon^{\prime})$
@@
22            $|\widetilde{M}_{i}(v)|=|M(v)|+\text{Lap}(2/\varepsilon^{\prime})$
24             $v$ releases $r_{v}$
25             for $u\in W_{r_{v}}(v)$ do
26                   $M(u)\leftarrow M(u)\cup\{v\}$
27                   $M(v)\leftarrow M(v)\cup\{u\}$
28
29
Algorithm 9 $\varepsilon$-LEDP Approximate Maximum $b$-Matching
We now restate and prove Theorem 1.4. See 1.4
The proof is implied by the following lemma which is a restatement with detailed constants.
Lemma 10.1.
For $\eta\in(0,1)$ and $b\geq\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}$, Algorithm 9 is $\varepsilon$-LEDP and outputs an implicit $b$-matching $M$ in the billboard model such that $|M|\geq\frac{1}{2}\operatorname{OPT}(b^{\prime})$ with probability at least $1-\frac{1}{n^{c}}$. Here $\operatorname{OPT}(b^{\prime})$ is the size of a maximum $b^{\prime}$-matching.
Proof.
For each iteration $i$, we will show that the number of nodes that $v=v_{i}$ matches with after iteration $i$ is either at least $b^{\prime}$, or every one of its unmatched neighbors is already matched to at least $b^{\prime}$ of their neighbors. Hence the final output satisfies the maximality condition of Lemma 4.2 and thus the approximation guarantees. We will show that this suffices to guarantee that our output is of comparable size to a maximal $b^{\prime}$-matching. Finally, we show that with high probability, we do not match with more than $b$ nodes.
First, all of the Laplace noise (Algorithms 9, 9, 9 and 9) are drawn from distributions with expectation 00. Let $X$ be a random variable drawn from one of these Laplace distributions. By Lemma A.1, we have that each of the following holds with probability $1-1/n^{3c}$:
Case I: Suppose that Algorithm 9 does not execute, i.e. $v$ has already satisfied its matching condition as evaluated on Algorithms 9, 9 and 9. Then we know that $|M(v)|+\nu_{i}(v)\geq\tilde{b}(v)$ by definition. We then have that
$|M(v)|$ $\geq b-{60}c\ln(n)/\varepsilon^{\prime}$
$\geq{b}-\frac{72c\ln(n)}{\varepsilon^{\prime}}$
$={b}-\frac{8\cdot 72c\ln(n)}{\eta\varepsilon}$ (by our choice of $\varepsilon^{\prime}$ in Algorithm 9)
$\geq{b}-\frac{576c\ln(n)}{\eta\varepsilon}$
$>{b^{\prime}}\,.$ (by $b\geq b^{\prime}+\frac{576c\ln(n)}{\eta\varepsilon}$)
In other words, $v$ was already matched to $b^{\prime}$ neighbors from a previous iteration.
Case II.a: Now, assume that Algorithm 9 does execute. Let $W_{0}(v)$ be the set of unmatched neighbors of $v$ which appear later in the ordering. First, consider the case that $|W_{0}(v)|<\frac{12c\ln(n)}{\eta^{2}}$. We know that either $|M(v)|<b^{\prime}$ or $|M(v)|\geq b^{\prime}$. If $|M(v)|\geq b^{\prime}$, then $v$ is already matched to sufficiently many neighbors and there is nothing to prove. Otherwise, conditioned on $\mathcal{E}$,
$|W_{0}(v)|+3\cdot\frac{12c\ln(n)}{\varepsilon^{\prime}}+{b^{\prime}}$ $\leq\frac{12c\ln(n)}{\eta^{2}}+\frac{36c\ln(n)}{\varepsilon^{\prime}}+{b^{\prime}}$
$\leq\frac{12c\ln(n)}{\eta^{2}}+\frac{216c\ln(n)}{\eta\varepsilon}+{b^{\prime}}$
$\leq\frac{238c\ln(n)}{\eta^{2}\varepsilon}+{b^{\prime}}$
$\leq b$
and all of $W_{0}(v)$ is matched.
Case II.b.1: For the remainder of the proof, we assume $k\coloneqq|W_{0}(v)|\geq\frac{12c\ln(n)}{\eta^{2}}$. As a reminder, we are in the case that Algorithm 9 does execute. Let $r^{*}$ be the $r$ chosen in Algorithm 9. If the returned $r^{*}>\log_{1+\eta}(k\eta^{2}/(12c\ln{n}))$, we know that the condition $|\widetilde{M}(v)|+|\widetilde{W}_{r}(v)|+12c\ln(n)/\varepsilon^{\prime}\leq b$ did not hold for $r=\log_{1+\eta}\left(\frac{k\eta^{2}}{12c\ln{n}}\right)$. Note here we use $k\geq\frac{12c\ln(n)}{\eta^{2}}$ so that $r\geq 0$ is well-defined. Thus, we have that $p_{r}=\frac{12c\ln(n)}{k\eta^{2}}$ so the expected size of $W_{r}(v)$ is $\frac{12c\ln(n)}{\eta^{2}}$; that is, $\mathbb{E}[|W_{r}(v)|]=\frac{12c\ln(n)}{\eta^{2}}$. By a multiplicative Chernoff Bound (Theorem A.2), the true size of $W_{r}(v)$ is at most $18c\log(n)/\eta^{2}$ with probability at least $1-\frac{1}{n^{4c}}$. Since the condition $|\widetilde{M}(v)|+|\widetilde{W}_{r}(v)|+12c\log(n)/\varepsilon^{\prime}\leq b$ didn’t hold for such $r$, we have that
$|M(v)|+|W_{r}(v)|+\frac{24c\ln(n)}{\varepsilon^{\prime}}$ $>b\,.$
In other words,
$|M(v)|$ $>b-|W_{r}(v)|-\frac{24c\ln(n)}{\varepsilon^{\prime}}$
$|M(v)|$ $>b-\frac{18c\ln(n)}{\eta^{2}}-\frac{24c\ln(n)}{\varepsilon^{\prime}}$ ( w.p. at least $1-\frac{1}{n^{4c}}$)
$|M(v)|$ $>b-\frac{18c\ln(n)}{\eta^{2}}-\frac{144c\ln(n)}{\eta\varepsilon}$ (by the choice of $\varepsilon^{\prime}$)
$|M(v)|$ $>b-\frac{162c\ln(n)}{\eta^{2}\varepsilon}$
$|M(v)|$ $>{b^{\prime}}\,.$ (since $b\geq{b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}}$)
In other words, $v$ was already matched from a previous iteration.
Case II.b.2: If $r^{*}\leq\log_{1+\eta}\left(\frac{k\eta^{2}}{12c\ln{n}}\right)$, there are two scenarios. Either $r^{*}=0$, then we match $v$ with all nodes which are still available so clearly the number of nodes $v$ is matched with is at least $b^{*}(v)$, or $r^{*}>1$. In the latter scenario, we first remark that $p_{r^{*}}\geq\frac{12c\ln(n)}{k\eta^{2}}$ so that
$\mathbb{E}[|W_{r^{*}-1}|]=(1+\eta)\mathbb{E}[|W_{r^{*}}|]\geq(1+\eta)\frac{12c\ln(n)}{\eta^{2}}\,.$
Thus we can apply a Chernoff bound on both $|W_{r^{*}-1}|,|W_{r^{*}}|$ and obtain high probability bounds. We also know that the threshold is exceeded for $r^{*}-1$, so we have $|\widetilde{M}(v)|+|\widetilde{W}_{r^{*}-1}(v)|+12c\ln(n)/\varepsilon^{\prime}\geq b$. Conditioning on $\mathcal{E}$, $|M(v)|+|{W}_{r^{*}-1}(v)|\geq b-36c\ln(n)/\varepsilon^{\prime}$. Thus by a Chernoff bound on $|W_{r^{*}-1}|$, we see that
$b-|M(v)|-\frac{36c\ln(n)}{\varepsilon^{\prime}}\leq|W_{r^{*}-1}|\leq(1+\eta)\mathbb{E}[|W_{r^{*}-1}|]=(1+\eta)^{2}\mathbb{E}[|W_{r^{*}}|]$
with probability $1-\frac{1}{n^{4c}}$. Now again by a Chernoff bound but now on $|W_{r^{*}}|$, we have
$|W_{r^{*}}|\geq\frac{(1-\eta)}{(1+\eta)^{2}}\left[b-|M(v)|-\frac{36c\ln(n)}{\varepsilon^{\prime}}\right]$
with probability $1-\frac{1}{n^{4c}}$.
Hence, if $|M(v)|<b^{\prime}$ (meaning it has not satisfied the matching criteria), it holds that
$|W_{r^{*}}(v)|$ $\geq\frac{(1-\eta)}{(1+\eta)^{2}}(b-|M(v)|-\frac{216c\ln(n)}{\eta\varepsilon})$
$\geq\frac{1-\eta}{(1+\eta)^{2}}\left(\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}-|M(v)|-\frac{216c\ln(n)}{\eta\varepsilon}\right)$
$\geq b^{\prime}-|M(v)|\,.$
It follows that $v$ is matched with at least $b^{\prime}$ neighbors at the end of this iteration.
Taking a union bound over all $r$ and $v\in V$, all Chernoff bounds hold regardless of the value of $r^{*}$ with probability at least $1-\frac{1}{n^{2c}}$ when $c\geq 3$. Taking into account the conditioning on $\mathcal{E}$, our algorithm succeeds with probability at least $1-\frac{1}{n^{c}}$ for $c\geq 3$.
In summary, at iteration $i$, we ensure that 1) either $v=v_{i}$ was already matched to at least $b^{\prime}$ neighbors during prior iterations, 2) or it will be matched with a sufficient number of descendents to satisfy a total of at least $b^{\prime}$ matched neighbors after iteration $i$, 3) or it will be matched to all descendents that do not yet have $b^{\prime}$ matched neighbors. The only property to check is whether there are any ancestors $u$ of $v$ that are not matched to $v$ but have yet to achieve $b^{\prime}$ matched neighbors in the last case. But this cannot happen since $u$ would have matched to all descendents that did not yet have $b^{\prime}$ matched neighbors during its iteration and $v$ in particular did not have $b^{\prime}$ matched neighbors.
$|M(u)|$ $<\tilde{b}(u)-\nu_{i}(u)$
$|M(u)|$ $<b-24c\log(n)/\varepsilon^{\prime}+24c\log(n)/\varepsilon^{\prime}$ (conditioning on $\mathcal{E}$)
$|M(u)|$ $<b.$
$=|M(v)|+|W_{r}(v)|.$
Thus in this case, $v$ is also matched to at most $b$ neighbors, as desired. ∎
We now restate and prove Theorem 1.5. See 1.5 The proof follows from the following corollary of Lemma 10.1.
Corollary 10.2 ($(2+\eta)$-Approximate Maximum $b$-Matching).
When $b\geq\frac{576c\ln(n)}{\eta^{3}\varepsilon}$ for a constant $\eta\in(0,\nicefrac{{1}}{{2}})$, Algorithm 9 is $\varepsilon$-LEDP and outputs an implicit $(2+24\eta)$-maximum $b$-matching in the billboard model, with probability at least $1-\frac{1}{n^{c}}$.
Before proceeding with the proof, we recall that $b$-matchings have the following integral relaxation [schrijver2003combinatorial, Theorem 33.2], where $\partial(U)$ denotes the set of outgoing edges from $U\subseteq V$.
$\sum_{uv\in E}x(uv)$ $\leq b$ $\forall v\in V$
$\sum_{e\in E[U]}x(e)+\sum_{f\in F}x(f)$ $\leq\lfloor\frac{1}{2}(b|U|+|F|)\rfloor$ $\forall U\subseteq V,F\subseteq\partial(U)$
$x$ $\in[0,1]^{E}$
Note that it suffices to consider constraints of the second form only when $b|U|+|F|$ is odd, but we include the redundant constraints for the sake of simplicity.
Proof.
Choose $b^{\prime}$ such that
$b=(1+12\eta)b^{\prime}\geq(1+11\eta)b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}\geq\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\frac{576c\ln(n)}{\eta^{2}\varepsilon}\,.$
Then we can apply Lemma 10.1 to deduce that Algorithm 9 outputs a $2$-approximate maximum $(\frac{b}{1+12\eta})$-matching.
Identify any $b$-matching $M$ as an indicator $\chi_{M}\in[0,1]^{E}$ of the $b$-matching polytope. Then $\frac{1}{1+12\eta}\chi_{M}$ is a feasible fractional solution of the $(\frac{b}{1+12\eta})$-matching polytope. Hence the size of a maximum $(\frac{b}{1+12\eta})$-matching is at least a $\frac{1}{1+12\eta}$ fraction of the size of a maximum $b$-matching. This concludes the proof. ∎
10.2 Node-Differentially Private $b$-Matchings
Similar to Section 6, we can implement our implicit $b$-matching algorithm in node-DP setting with the help of an arboricity sparsifier. However, this requires extending Solomon’s sparsifier [solomon2021local] to the setting of $b$-matchings. Doing so and repeating the steps in Section 6 yields a proof of Theorem 1.9. See 1.9
The rest of this section serves to prove the necessary sparsifier extension.
10.2.1 Bounded Arboricity Sparsifiers for $b$-Matchings
In order to extend the $b$-matching guarantees to the node-DP setting, we prove an extension of Solomon’s sparsifier (Theorem 6.2) to the case of $b$-matchings with an $O(b)$ additive increase in the degree threshold. This may be of independent interest beyond node differentially private matchings.